OpenStudy (anonymous):

Choose the correct simplification of the expression 4/y^-3 1. 4y^3 2. 4/y^3 3. y^3/4 4.This expression is already simplified.

OpenStudy (anonymous):

It would be 1.

OpenStudy (anonymous):

\[x^{-n} = 1/x^n\]

OpenStudy (anonymous):

so you get: \[\frac{1}{\frac{1}{y^3}} = y^3\]

OpenStudy (anonymous):

so: \[4y^3\]